scott is 4 times olderthan bethany. In 6 years scott will only be 3 times older than bethany, what are 2 equations? and how old are they? please help I am totally stuck...thanks heaps
Answer by stanbon(75887)

scott is 4 times olderthan bethany. In 6 years scott will only be 3 times older than bethany, what are 2 equations?
------
Equations:
s = 4b
(s+6) = 3(b+6)
------------------------
Substitute for "s" and solve for "b":
4b + 6 = 3b + 18
b = 12 (Bethany's age now)
----
Solve for "s":
s = 4b
s = 4*12 = 48 (Scott's age now)
===================
Cheers,
Stan H.
==============